OpenStudy (anonymous):

Can you explain this word problem? (i am not good at word problems. There are 110 calories per 28.4 grams of cereal X. Find how many calories are in 45.44 grams of this cereal.

OpenStudy (anonymous):

its 176 calories when you divide 110 by 28.4 you get no.of calories for each gram then you multiply it with 45.44 then you will get the answer or 110/28.4 = x/45.44 you will get x=176

OpenStudy (anonymous):

176 calories